OpenStudy (anonymous):

If xy+z=y what is x in terms of y and z? a) y+z/y b) y-z/z c) y-z/y d) 1-z e) z-y/y

hartnn (hartnn):

to isolate x first subtract z from both sides, what do u get ?

OpenStudy (anonymous):

xy=y-z i think

hartnn (hartnn):

yup, now since y is 'multiplied' by x on left, to get only x, u divide both sides by y, what u get?

OpenStudy (anonymous):

x=-z

hartnn (hartnn):

nopes, u divide it like this : \(\huge\frac{xy}{y}=\frac{y-z}{y}\) to get \(\huge x=\frac{y-z}{y}\) got it ?

OpenStudy (anonymous):

yes than you

hartnn (hartnn):

welcome :)